OpenStudy (anonymous):

There is a rectangle where the length is double the width and the perimeter is 24 feet. Write an equation that can be used to determine the length and width of the rectangle, and use it to find the dimensions. Indicate the step where variables are defined, the equation used to solve, and the final answer.

OpenStudy (anonymous):

helpp

ganeshie8 (ganeshie8):

say, length = l, width = w

ganeshie8 (ganeshie8):

try to form equations using the given info

ganeshie8 (ganeshie8):

length is double the width \(l = 2w\) -----------------------------(1)

ganeshie8 (ganeshie8):

thats ur first equation

OpenStudy (anonymous):

2w + 4l = 24??

ganeshie8 (ganeshie8):

close

ganeshie8 (ganeshie8):

perimeter is 24 feet \(2(l+w) = 24\) \(l+w = 12\)----------------------------(2)

ganeshie8 (ganeshie8):

your second equation

ganeshie8 (ganeshie8):

using those two equations, you need to find the length \(l\), and width \(w\)

OpenStudy (anonymous):

so the equation is 2w + l + l = 24

OpenStudy (anonymous):

im confused

ganeshie8 (ganeshie8):

both equations you need to use

ganeshie8 (ganeshie8):

let me show you how to solve

ganeshie8 (ganeshie8):

take equation (2) \(l + w = 12\) plugin \(l = 2w\)

ganeshie8 (ganeshie8):

\(2w + w = 12\)

OpenStudy (anonymous):

so 4 is the answer

ganeshie8 (ganeshie8):

\(3w = 12\) \(w = 4\)

ganeshie8 (ganeshie8):

width is 4, you still need to find the length \(l\)

ganeshie8 (ganeshie8):

for \(l\), just plugin w=4, in equation (1)

ganeshie8 (ganeshie8):

\(l = 2w\) = 2(4) = 8

ganeshie8 (ganeshie8):

so, the dimensions of rectangle are : length = 8 width = 4

ganeshie8 (ganeshie8):

see if that makes sense. read ur question completely. it wants you to explain how we worked out... so spend some time on this ok :)